when two numbers are added together, the result is 45. twice the greater number is six more than five times the lesser number
Answer by mando271(42)

we'll let the numbers be x and y with x being the lesser
x+y=45
2y=5x+6
now we'll use substitution method to solve. I'll use the 1st equation to solve for y.
x+y=45
y=45-x
now plug this into the 2nd equation
2(45-x)=5x+6
90-2x=5x+6
-7x=-84
x=12
now plug this into our 1st solution to find y
y=45-12
y=33
so, x=12 and y=33